# Quantum noise modeling through Reinforcement Learning

### v2

@@ I Introduction
One important unsolved technological question regards the practical applicability of Noisy Intermediate Scale Quantum (NISQ) [1] computers. Despite the expectation that quantum computers will outperform classical computers in certain computational tasks [2, 3, 4], the usability and reliability of NISQ devices are hindered by a large error rate. These errors arise from gate infidelities, unwanted environmental interactions, thermal relaxation, measurement errors, and cross-talk [5, 6, 7, 8, 9]. Currently, there are techniques to mitigate these errors [10, 11, 12, 13, 14]. However, it has been demonstrated that any quantum circuit for which error mitigation is efficient must be classically simulable [15]. Therefore, it is widely regarded that quantum advantage will only be achieved with future generations of fault-tolerant quantum devices [16, 17, 18, 19, 20, 21].
Despite the imperfect results obtained on NISQ devices, numerous algorithms have been developed and implemented on this hardware. Machine learning-inspired models, in particular, have shown promising results in recent years [22, 23, 24, 25, 26, 27, 28, 29]. Testing and developing new algorithms in the NISQ era can be challenging due to limited access to quantum chips. Furthermore, the currently accessible quantum computers in the cloud are highly demanded, resulting in lengthy waiting queues [30]. In this context, emulating the noise of these devices emerges as an alternative to accelerate circuit testing.
@@
This work aims to develop a model capable of learning hardware-specific noise for use in circuit simulations. This goal is further motivated by the limited availability of noise modeling or noise prediction techniques [31, 32, 33, 34]. Our approach employs Reinforcement Learning (RL) [35, 36, 37] to train an agent to add noise channels that replicate the noise pattern of a specific quantum chip. This method minimizes heuristic assumptions about the noise model, thereby enhancing the adaptability and generalization properties. The algorithm has been validated on both simulations and real quantum devices hosted at the Quantum Research Center [38] of the Technology Innovation Institute (TII) in Abu Dhabi, demonstrating its ability to accurately predict different noise patterns.
The Qibo framework [39, 40] was used for the realization of this work. It provides Qibo [39, 40] as a high-level language API for crafting quantum computing algorithms, Qibolab [41, 42] as a tool for quantum control, and Qibocal [43] for conducting quantum characterization and calibration routines.
All the code developed for this work is available on GitHub  [^1] https://github.com/qiboteam/rl-noisemodel. It is possible to use the code to reproduce the results as well as testing the algorithm under different noise conditions. The code is intended to be easily customizable in order to allow users to define their own RL agents for specific applications.
@@
The outline is as follows. Section II introduces the basic concepts necessary to understand the proposed algorithm, noise in quantum circuits and reinforcement learning. Section III provides a detailed description of the reinforcement learning algorithm, focusing on its training and use for noise prediction. Section IV presents the results obtained with the proposed algorithm on both simulations and real quantum devices, and compares its performance with other noise predictors. Section V demonstrates examples of the algorithm’s use-cases for famous quantum algorithms.
@@ II Background
This section provides an overview of the essential concepts required to understand the proposed noise simulation algorithm. Specifically, we briefly discuss noise in quantum circuits and introduce reinforcement learning.
II.1 Noise in the quantum circuit model
@@
It is common to make certain assumptions about the inherent errors. One typical assumption is to distinguish the ideal gates from the errors, considering them as separate processes. It is also often assumed that these errors break down into spatially uncorrelated errors that affect the idle qubits, and an average error that is operation-independent, impacting only the qubits that are being manipulated. Within this context, it is useful to classify the various types of noise introduced above into two groups: coherent and incoherent. Coherent noise, typically resulting from minor miscalibrations in control parameters, tends to produce similar errors in successive executions of a quantum circuit, thereby introducing a systematic bias in the output. It is important to note that coherent noise preserves the purity of the state and, once identified, can be corrected [46, 47, 48]. On the other hand, incoherent noise can be viewed as processes that cause entanglement between the system and its environment.
The errors and imperfect operations are typically represented using the formalism of quantum channels, i.e., trace-preserving completely positive maps of density matrices into density matrices. In the ideal operation of a quantum computer, a positive map can be just a unitary transformation $\varepsilon(\rho)=U\rho U^{\dagger}$, where $U$ describes a quantum gate. Coherent noise is unitary, and we model it using single qubit rotation gates ($R_{x}$, $R_{y}$, $R_{z}$). For instance, a coherent error could introduce an unintended deviation $\epsilon$ in the $x$ direction during the application of $R_{j}(\theta)$, altering the state $\rho=R_{j}(\theta)\rho_{0}R_{j}(\theta)^{\dagger}$ to
@@
where $A_{1}=\ket{0}\bra{0}+\sqrt{1-\gamma}\ket{1}\bra{1}$ | and $A_{2}=\sqrt{\gamma}\ket{0}\bra{1}$ |, and $\gamma$ represents the decay probability from |$\ket{1}$ to |$\ket{0}$. While it is possible to add more complexity to the modeling of incoherent noise, we aim for an effective description with few parameters to prevent overfitting. Preliminary tests showed no significant performance enhancement when adding other channels.
@@
Multiple noise channels can be combined in order to construct complex noise models [33, 49]. The parameters of these channels can be inferred either from calibration results or directly from the execution of quantum circuits, as demonstrated in this work.
For a simplified noise modeling, we can employ a technique known as Randomized Benchmarking (RB) [50, 51, 52]. RB enables us to efficiently estimate the average error magnitude across a set of quantum gates, with resource requirements scaling polynomially with the number of qubits. A RB protocol employs sequences of varying lengths of randomly chosen $n$-qubit Clifford gates, where the ideal composite operation of the sequence is the identity. To produce such a sequence of depth $m+1$, each of the first $m$ gates in the sequence are picked randomly from the Clifford group. The last gate is then uniquely determined as the Clifford element which inverts the composition of the previous $m$ gates. In the presence of noise, the actual sequence of Clifford gates does not represent an identity operation. Instead, there is a certain survival probability of measuring the initial state after performing the sequence. By estimating this survival probability across multiple independent random sequences of increasing depth, we can fit an exponential decay of the form
$P_{l}=ap^{l}+b\,.$ (4)
The average error map, therefore, behaves like a completely depolarizing model with an average depolarizing parameter $p$. This parameter can be used to estimate the average gate fidelity of $n$-qubit quantum gates. Employing RB as a noise predictor involves extracting the decay parameter and introducing a depolarizing channel with a depolarizing parameter equal to the decay parameter after each gate. We employ this technique to establish a depolarization noise model, serving as a benchmark for our algorithm. Although this model, which projects all noise sources onto the depolarizing channel, may seem unrealistic, it provides a useful baseline for comparison.
II.2 Reinforcement learning
@@
Many optimization methods have been developed in recent years to improve reinforcement learning convergence and stability during training [56]. In our work, we have obtained the best results using the Proximal Policy Optimization (PPO) [57]. This optimization method provides increased stability by constraining the policy update to a proximity region. Moreover, PPO can handle a continuous action space, which is necessary to accurately model quantum noise. One drawback is its sensitivity to hyperparameters, as poorly chosen values can impact convergence and overall performance.
@@ III.1 Circuit Representation
To train the RL agent, we need to represent a quantum circuit as an array that can be readily processed by the policy neural network. In the following we refer to this array as the Quantum Circuit Representation (QCR). The QCR has a shape of $[\,qubits,\>depth,\>encoding\,]$. The first dimension corresponds to the circuit’s qubits, while the second dimension represents the circuit’s moments, i.e., a collection of gates that can be executed in parallel. The $encoding$ dimension encodes the information regarding gates and noise channels acting on a specific qubit at a specific circuit moment. Its dimension is determined by the total number of native gates and noise channels the model allows for.
In detail, considering a set of $n$ single-qubit native gates ${G_{i}}_{i=1}^{n}$, the initial $n$ entries of the $encoding$ will contain a one-hot encoding of our native set. Namely, in the presence of the gate $G_{i}$, all the entries are set to 0 except for the $i$-th entry that is set to 1. Specifically, in our case, we use $n=2$ to encode the presence of an $R_{x}$ or $R_{z}$ gate. The $n+1$ entry is set to $1$ if a two-qubit gate is present and 00 otherwise, with the $CZ$ gate being the native two-qubit gate in this work. For single-qubit circuits, this entry is always zero. The $n+2$ entry encodes the rotation angle of single-qubit gates, normalized to the range $[0,1]$. The remaining entries detail the parameters of the noise channels in the following sequence: Depolarizing channel, Amplitude damping channel, $R_{z}$ coherent error, and $R_{x}$ coherent error. If a noise channel is absent, the corresponding parameter value is set to zero. An example of the QCR for a two-qubit circuit, obtained using this procedure, is illustrated in Figure 1.
{quantikz}
[…]
$R_{z}$ $R_{x}$ $CZ$ $\theta$ $\operatorname{Dep}$ $\operatorname{Dam}$ $\operatorname{Coh_{z}}$ $\operatorname{Coh_{x}}$
@@
Figure 1: Example of a two-qubit quantum circuit (top) and its vector representation (bottom).
To enable the agent to adapt to circuits of varying depths, we further introduce the concept of kernel size $k$, similar to the kernels used in convolutional neural networks (CNN) [61]. The kernel size establishes a “window” that restricts the number of circuit moments the agent can observe at any given time. For instance, with $k=3$, the agent only observes the current moment and the immediately preceding and following ones. The window’s center starts from the first moment and slides one position at each step until the circuit’s end is reached. This approach is based on the heuristic assumption that a gate’s noise is most influenced by its temporally proximate gates. At a given moment $m$, therefore, a window $(m-\frac{k-1}{2},\,m+\frac{k-1}{2})$ is extracted from the complete QCR of the circuit, effectively yielding a fixed dimension $[\,qubits,\>k,\>encoding\,]$ tensor that is fed as input to the agent.

### v3

@@ I Introduction
One important unsolved technological question regards the practical applicability of Noisy Intermediate Scale Quantum (NISQ) [66] computers. Despite the expectation that quantum computers will outperform classical computers in certain computational tasks [81, 32, 96], the usability and reliability of NISQ devices are hindered by a large error rate. These errors arise from gate infidelities, unwanted environmental interactions, thermal relaxation, measurement errors, and cross-talk [47, 46, 95, 36, 64]. Currently, there are techniques to mitigate these errors [56, 30, 87, 83, 84, 93]. However, it has been demonstrated that any quantum circuit for which error mitigation is efficient must be classically simulable [80]. Therefore, it is widely regarded that quantum advantage will only be achieved with future generations of fault-tolerant quantum devices [49, 29, 28, 89, 82, 9]. Despite the imperfect results obtained on NISQ devices, numerous algorithms have been developed and implemented on this hardware. Machine learning-inspired models, in particular, have shown promising results in recent years [3, 16, 12, 25, 5, 74, 75, 19].Testing and developing new algorithms in the NISQ era can be challenging due to limited access to quantum chips. Furthermore, the currently accessible quantum computers in the cloud are highly demanded, resulting in lengthy waiting queues [72]. In this context, emulating the noise of these devices emerges as an alternative to accelerate circuit testing.
@@
This work aims to develop a model capable of learning hardware-specific noise for use in circuit simulations. Our approach employs Reinforcement Learning (RL) [86, 44, 90] to train an agent to add noise channels that replicate the noise pattern of a specific quantum chip. This method minimizes heuristic assumptions about the noise model, thereby enhancing the adaptability and generalization properties. The algorithm has been validated on both simulations and real quantum devices hosted at the Quantum Research Center [69] of the Technology Innovation Institute (TII) in Abu Dhabi, demonstrating its ability to accurately predict different noise patterns.
@@
The outline is as follows. Section II introduces the basic concepts necessary to understand the proposed algorithm, noise in quantum circuits and reinforcement learning. Moreover, we review previous works on similar topics to provide more context of the work in relation to recent applications of reinforcement learning in quantum computing and noise simulation techniques. Section III provides a detailed description of the reinforcement learning algorithm, focusing on its training and use for noise prediction. Section IV presents the results obtained with the proposed algorithm on both simulations and real quantum devices, and compares its performance with other noise predictors. Section V demonstrates examples of the algorithm’s use-cases for famous quantum algorithms.
The Qibo framework [24, 22] was used for the realization of this work. It provides Qibo [24, 22] as a high-level language API for crafting quantum computing algorithms, Qibolab [23, 65] as a tool for quantum control, and Qibocal [62] for conducting quantum characterization and calibration routines. All the code developed for this work is available on GitHub  [^1] https://github.com/qiboteam/rl-noisemodel. It is possible to use the code to reproduce the results as well as testing the algorithm under different noise conditions. The code is intended to be easily customizable in order to allow users to define their own RL agents for specific applications.
@@ II Background
This section provides an overview of the essential concepts required to understand the proposed noise simulation algorithm. Specifically, we briefly discuss noise in quantum circuits, the randomized benchmarking technique and introduce reinforcement learning. We also review some related results obtained in the context of noise simulation and reinforcement learning applied to quantum computing.
II.1 Related works
The characterization and modeling of quantum noise is an important research area that has yielded significant results in recent years [92, 31, 34]. While standard noise tomography techniques such as randomized benchmarking are useful for obtaining a standardized evaluation of noise, they lack the ability to capture specific characteristics that are essential for realistic simulations. These techniques can be improved by incorporating non-Markovian noise and realistic finite-time gate models, as proposed by Brillant et al. [8]. Other studies focus on reducing the resources required to perform these techniques and on maximizing the information extracted from experimental data [60, 41, 14].Another line of research involves developing efficient algorithms to fit a Lindbladian to tomography data [52, 60, 88], providing compact noise models and quantitative measures of non-Markovianity. Finally, it is worth mentioning the tensor network approach proposed by Mangini et al. [53], which enables the study of correlated noise in systems with up to 20 qubits.Various techniques based on machine learning have also been proposed for noise characterization. For instance, Zlokapa et al. introduced a deep learning method to identify noise patterns, aiming to reduce noise levels during the transpilation process [97]. Canonici et al. trained machine learning models on measurement data to infer noise parameters in neutral-atom quantum processors and additionally employed a reinforcement learning agent to design corrective control pulses that mitigate measurement noise [10].
Separately, reinforcement learning has been explored for a variety of quantum control and calibration tasks. In quantum gate design, several studies apply deep RL to optimize pulse shapes or gate parameters. For instance, Niu et al. trained an RL agent to optimize two-qubit gate pulses in the presence of leakage and stochastic errors [58]. In superconducting systems, Baum et al. experimentally employed deep RL to autonomously design single- and two-qubit gates without requiring knowledge of the device Hamiltonian [2].Reinforcement learning has also been applied to improve qubit readout [40]. Chatterjee et al. demonstrated on superconducting hardware that an RL agent can tailor readout pulses to achieve state-of-the-art fidelity while significantly reducing readout time [13]. Similar applications of RL methods have been proposed for automated calibration and tuning. For example, Crosta et al. [18] developed a model-free RL agent that periodically adjusts calibration parameters to maintain optimal performance, without relying on an explicit noise model of the environment. Related frameworks have been introduced by K. Young [94] and Nguyen et al. [55] to accelerate the calibration of superconducting logic gates.Finally, several works have applied RL to quantum error correction and quantum algorithm design. Nautrup et al. used RL to adapt surface-code error-correction circuits to unknown noise channels [57]. Olle et al. extended this approach to discover new error correction codes and encoding circuits tailored to a given hardware’s gate set, connectivity, and noise model [59].
II.2 Noise in the quantum circuit model
@@
It is common to make certain assumptions about the inherent errors. One typical assumption is to distinguish the ideal gates from the errors, considering them as separate processes. It is also often assumed that these errors break down into spatially uncorrelated errors that affect the idle qubits, and an average error that is operation-dependent, impacting only the qubits that are being manipulated. Within this context, it is useful to classify the various types of noise introduced above into two groups: coherent and incoherent. Coherent noise, typically resulting from minor miscalculations in control parameters, tends to produce similar errors in successive executions of a quantum circuit, thereby introducing a systematic bias in the output. It is important to note that coherent noise preserves the purity of the state and, once identified, can be corrected [11, 45, 6]. On the other hand, incoherent noise can be viewed as processes that cause entanglement between the system and its environment.
The errors and imperfect operations are typically represented using the formalism of quantum channels, i.e., trace-preserving completely positive maps of density matrices into density matrices. In the ideal operation of a quantum computer, a positive map can be just a unitary transformation $\varepsilon(\rho)=U\rho U^{\dagger}$, where $U$ describes a quantum gate. Coherent noise is unitary, and we model it using single-qubit rotation gates ($R_{x}$, $R_{y}$, $R_{z}$). For instance, a coherent error could introduce an unintended deviation $\epsilon$ in the $x$ direction during the application of $R_{j}(\theta)$, altering the state $\rho=R_{j}(\theta)\rho_{0}R_{j}(\theta)^{\dagger}$ to
@@
[…]
@@
Multiple noise channels can be combined in order to construct complex noise models [31, 4]. The parameters of these channels are usually inferred from calibration results. In this work, we show that it is possible to use ML to fit a noise model directly from the execution of quantum circuits.
II.3 Randomized benchmarking
Randomized benchmarking (RB) is widely recognized as one of the most effective experimental techniques for evaluating the quality of quantum operations in quantum computing platforms [27, 38, 37, 35]. Its widespread adoption is due to practical advantages, particularly its polynomial scalability with the number of qubits and its robustness to state preparation and measurement errors.
At the core of the RB protocol lies a gateset $G$, typically chosen to be a finite group of quantum gates such as the Clifford group [51]. The primary objective of RB is to estimate the average gate fidelity of this gateset, providing a reliable measure of the performance of the underlying quantum device.
The protocol proceeds by randomly sampling a sequence of $m$ gates from the gateset $G$, applying them to an initial quantum state $\rho$, and then applying an additional gate designed to invert the net effect of the sequence in the absence of noise. This last gate ensures that the final state should ideally return to $\rho$ if all gates were implemented perfectly. The actual final state is then measured to estimate the survival probability $p_{m}$, defined as the probability of finding the system in the initial state after the noisy evolution. This process is repeated over many random sequences of the same length $m$, and the resulting probabilities are averaged to obtain a single estimate of $p_{m}$. By varying the sequence length $m$, one obtains a decay curve, which reflects the accumulation of errors and allows extraction of the average fidelity of the gateset.
The Clifford group is typically chosen as the gateset $G$ in RB because it enables a simple model for how errors accumulate. Under the assumption of gate-independent, completely positive and trace-preserving (CPTP) noise, the average survival probabilities $\{p_{m}\}$ can be well-approximated by a single exponential decay of the form
$p_{m}=af^{m}+b\,$, (4)
where $a$ and $b$ depend on state preparation and measurement errors, and the decay parameter $f$ reflects the average fidelity of the implemented gates [51]. This behavior arises because averaging over the Clifford group effectively transforms arbitrary noise into a depolarizing channel, which accumulates in an exponential decay with the sequence length $m$. This property, unique to unitary 2-designs like the Clifford group, makes it especially well-suited for benchmarking purposes  [^2] In RB protocols that employ non-Clifford gatesets, the standard exponential decay model (4) no longer holds. Instead, the decay behavior must be described by a generalized multi-exponential form $p_{m}=\sum_{\lambda\in J_{G}}A_{\lambda}f_{\lambda}^{m}$, where $J_{G}$ is an index set determined by the structure of the gateset, the $f_{\lambda}$ are gateset-dependent quality parameters, and the coefficients $A_{\lambda}$ depend only on measurement errors [37]..
Randomized benchmarking can be repurposed as a simplified noise modeling tool by treating the average error channel as an effective depolarizing channel with depolarizing rate 1f1-f. In this approach, a depolarizing channel with depolarizing rate 1f1-f is inserted after each gate to emulate the observed error behavior (see Appendix A). We adopt this technique to construct a baseline depolarizing noise model, which serves as a reference for evaluating our algorithm. While this model simplifies the noise landscape by projecting all errors onto a single depolarizing channel, potentially overlooking coherent or gate-dependent contributions, it remains a valuable benchmark for comparison. Importantly, we do not sample gates randomly from the Clifford group; instead, we work with a fixed set of Clifford gates native to the quantum hardware under study. In this setting, the assumptions underlying standard RB theory may not hold, and the noise may not fully depolarize. Consequently, the RB decay may deviate from a pure exponential form.
II.4 Reinforcement learning
@@
Many optimization methods have been developed in recent years to improve reinforcement learning convergence and stability during training [61]. In our work, we have obtained the best results using the Proximal Policy Optimization (PPO) [79]. The idea behind PPO lies in restricting the policy updates to avoid too large changes in the policy, which may lead to instability during training and, as people say in jargon, to “fall off the cliff”. Namely, too large updates may lead the agent to regions with low expected cumulative reward, that may take a very long time to escape from. This is implemented in practice by clipping the loss function to discourage big policy updates:
$\mathcal{L}_{t}(\theta)=\hat{E}_{t}[\min(r_{t}(\theta)\cdot\hat{A}_{t},\rm{clip}(r_{t}(\theta),\epsilon)\cdot\hat{A}_{t})]\;.$ (5)
The probability ratio $r_{t}(\theta)$ is defined as:
$r_{t}(\theta)=\frac{\pi^{t}_{\theta}(a_{t}|s_{t})}{\pi^{t-1}_{\theta}(a_{t}|s_{t})}\;.$ (6)
It evaluates the change in the policy going from step t1t-1 to $t$, and is limited in the final loss through the clip function:
$\rm{clip}(r_{t}(\theta),\epsilon)=\begin{cases}1+\epsilon&\text{if }r_{t}(\theta)>1+\epsilon,\\ 1-\epsilon&\text{if }r_{t}(\theta)<1-\epsilon,\\ r_{t}(\theta)&\rm{elsewhere}.\end{cases}$ (7)
This way big changes in the policy will still have a limited impact on the loss. This clipped ratio is then multiplied by the advantage factor $\hat{A}_{t}$, as in the non-clipped case, which measures how good is on average the action taken at step t over all the other possible actions. Finally, the minimum between the clipped and non-clipped reward is taken to provide “pessimistic predictions” and, thus, discourage larger jumps in the policy. PPO often outperforms algorithms like Q-learning, in environments with continuous action spaces or partial observability, where Q-learning struggles due to the need for discretization. Compared to Soft Actor-Critic (SAC), which may achieve better asymptotic performance in highly stochastic settings due to its entropy regularization, we preferred PPO for its computational simplicity and lower variance in updates, while yielding satisfactory results.
@@ III.1 Circuit Representation
To train the RL agent, we need to represent a quantum circuit as an array that can be readily processed by the policy neural network; we refer to this array as the Quantum Circuit Representation (QCR). In the following, we refer to a circuit moment as a collection of gates that act on a non-overlapping set of qubits and are conceptually executed in parallel. The circuit depth is defined as the total number of circuit moments, assuming that the division of the circuit into moments is optimized to minimize their number. The QCR has a shape of [qubits,depth,encoding][\,qubits,\>depth,\>encoding\,]. The first dimension corresponds to the circuit’s qubits, while the second dimension represents the circuit’s moments. The encodingencoding dimension encodes the information regarding gates and noise channels acting on a specific qubit at a specific circuit moment. Its dimension is determined by the total number of native gates and noise channels the model allows for.
In detail, considering a set of $n$ single-qubit native gates $\{G_{i}\}$, the initial $n$ entries of the encodingencoding will contain a one-hot encoding of our native set. Namely, in the presence of the gate $G_{i}$, all the entries are set to 0 except for the $i$-th entry that is set to 1. In the presence of a two-qubit gate, $CZ$ in our case, the entry relative to both the target and control qubit is set to 1. Specifically, in our case, we use $n=2$ to encode the presence of an $R_{x}$ or $R_{z}$ gate. The $n+1$ entry is set to $1$ if a two-qubit gate is present and 0 otherwise, with the $CZ$ gate being the native two-qubit gate in this work. For single-qubit circuits, this entry is always zero. The $n+2$ entry encodes the rotation angle of single-qubit gates, normalized to the range $[0,1]$. The remaining entries detail the parameters of the noise channels in the following sequence: Depolarizing channel, Amplitude damping channel, $R_{z}$ coherent error, and $R_{x}$ coherent error. If a noise channel is absent, the corresponding parameter value is set to zero. An example of the QCR for a two-qubit circuit, obtained using this procedure, is illustrated in Figure 1.
                               $q_{1}$ $R_{x}(\pi)$ $R_{x}(\pi)$ $\operatorname{Dep}(0.1)$ $\operatorname{Dep}(0.1)$ $q_{2}$ $\operatorname{Coh_{x}}(0.1)$ $\operatorname{Coh_{x}}(0.1)$ $\operatorname{Dam}(0.2)$ $\operatorname{Dam}(0.2)$ $R_{x}(\pi/2)$ $R_{x}(\pi/2)$ $\operatorname{Coh_{z}}(0.05)$ $\operatorname{Coh_{z}}(0.05)$ Moment 1 Moment 2 Moment 3
$R_{z}$ $R_{x}$ CZCZ $\theta$ $\operatorname{Dep}$ $\operatorname{Dam}$ $\operatorname{Coh_{z}}$ $\operatorname{Coh_{x}}$
@@
Figure 1: Example of a two-qubit quantum circuit (top) and its vector representation (bottom). A two-qubit circuit of depth three is represented as a tensor of size (2, 3, 8), where the first entry identifies the qubit, the second denotes the circuit moment, and the third specifies the type of gate or noise channel.
To enable the agent to adapt to circuits of varying depths, we further introduce the concept of kernel size $k$, similar to the kernels used in convolutional neural networks (CNN) [33]. The kernel size establishes a “window” that restricts the number of circuit moments the agent can observe at any given time. For instance, with $k=3$, the agent only observes the current moment and the immediately preceding and following ones. The window’s center starts from the first moment and slides one position at each step until the circuit’s end is reached. To ensure that the first and last moments of the circuit are at the center of a full-sized window, we apply a symmetric zero-padding of length (k1)/2(k-1)/2 at the edges of the QCR. This approach is based on the heuristic assumption that a gate’s noise is most influenced by its temporally proximate gates. At a given moment $m$, therefore, a window $(m-\frac{k-1}{2},\,m+\frac{k-1}{2})$ is extracted from the complete QCR of the circuit, effectively yielding a fixed dimension [qubits,k,encoding][\,qubits,\>k,\>encoding\,] tensor that is fed as input to the agent.